# On the Use of Large Language Models in Mutation Testing

### v2

@@ 3.2. Datasets
Table 1. Real Bugs Used in Our Experiment
@@
We intend to evaluate our approach with real bugs, and thus we need to use bug datasets with the following properties:
@@
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023a; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total. However, from Table 2 and Table 1, we observe the time spans of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder [^1] https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
3.3. Mutation Generation via LLMs
Given the location of a Java program, we extract essential information (e.g., the context method or the corresponding unit tests) to formulate prompts instructing mutation generation, then feed prompts to the selected LLM. Once the LLM returns its responses, we filter out mutations that are non-compilable, redundant, or identical to the original code, and return the generated mutations. As shown in Section 3.4, our study supports a comprehensive set of metrics for evaluating mutations.
3.3.1. Studied Models
We aim to comprehensively compare the existing LLMs in their ability to generate mutations from various perspectives. This includes evaluating the capabilities of models that are fine-tuned on different base models and contrasting the performance of commercial closed-source models with open-source alternatives. In our study, we select a representative set of LLMs with code capabilities. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from GPT family, including GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-Mini (Achiam et al., 2023). Moreover, we employ the popular open-sourced LLMs that are used by recent studies (Ni et al., [n. d.]; Xia et al., 2024; Ma et al., 2024; Yang et al., 2024), including DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023). The details of these models are shown as Table 2. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.
Table 2. Studied LLMs
Model Type Studied Model Base Model Training Data Time Release Time Size
Closed GPT-3.5-Turbo GPT 2021/09 2023/03
GPT-4o GPT 2023/10 2024/05
GPT-4o-Mini GPT 2023/10 2024/07
Open StarChat-$\beta$-16b StarCoder 2023/06 16B
CodeLlama-Instruct-13b Llama 2023/08 13B
DeepSeek-Coder-V2-236b DeepSeek 2023/09 2024/07 236B
3.3.2. Prompts
Prompts are crucial for LLMs because they guide the models’ responses and determine the direction of the generated output. A well-crafted prompt can significantly influence the utility, relevance, and accuracy of the code produced by an LLM. This subsection introduces how we design the prompts.
@@ 3.3.2. Prompts
Figure 1 shows the default prompt template in our study. To design effective prompts for mutation generation, we follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments.
In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs. This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
@@
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our default prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.

### v3

@@ 2.3. Datasets
Table 2. Real-world Bugs Used in Our Study
@@
As one of the major tasks of our study is to evaluate the behavioral similarity between mutations and real-world bugs, we selected datasets that include real-world bugs along with their corresponding fault-revealing test cases. Specifically, we employ the datasets with the following properties:
@@
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total. However, from Table 1 and Table 2, we observe the periods of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder [^2] https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
2.4. Evaluation Metrics
In this subsection, we present the evaluation metrics used in our study to assess the quality of the generated mutations comprehensively. We evaluate the mutations from multiple perspectives as before mentioned. Note that only relying on the mutation score is insufficient for a meaningful assessment. A higher mutation score does not necessarily indicate better quality mutations, as it only reflects the proportion of killed mutants rather than their effectiveness. However, we still report the mutation score of each mutation generation approach later in Section 3.
To better understand the metrics, we first illustrate the categories of mutations. Let all generated mutations be the set $A$, where LLMs may generate syntactically incorrect mutations that can not pass compilation. Thus we denote the compilable mutations as the set $C$. Note that the set $C$ contains two types of mutations that should be removed, the duplicate mutations (denoted as $D$) and equivalent mutations (denoted as $E$). Therefore, we have $(C\supseteq D)\wedge(C\supseteq E)\wedge(D\cap E=\emptyset)$. The duplicate mutations refer to those identical to the original code or other mutations. In mutation testing, the ideal set of mutations to employ is from the set $(C-E-D)$, excluding equivalent mutations that have no contributions to discovering the weakness of tests. However, determining the equivalent mutations set $E$ is undecidable, so we follow the common practice and use the set $(C-D)$ as a practical approximation to perform mutation testing (Diniz et al., 2021; Tian et al., 2022; Zhang et al., 2022; Wang et al., 2025).
2.4.1. RQ1: Effectiveness
Evaluating the effectiveness of mutation generation approaches requires assessing how well the generated mutations mimic real bugs. A key aspect of this evaluation is understanding whether LLM-generated mutations exhibit fault characteristics that align with real-world defects.
To quantify this effectiveness, we employ multiple metrics that capture different aspects of behavioral similarity between mutations and real bugs. These metrics assess whether bug-triggering tests can detect LLM-generated mutations (i.e., RQ1.1), how closely mutations resemble real faults based on test case failures (i.e., RQ1.2), and the diversity of syntactic changes introduced by different mutation generation techniques (i.e., RQ1.3). The metrics of these sub-RQs are listed as bellow.
RQ1.1: Real Bug Detection Rate:  Following the approach of Just et al. (Just et al., 2014b), we evaluate the effectiveness of mutation generation by measuring its ability to detect real bugs using developer-written tests. For a given real bug (which may span multiple lines), we apply a mutation generation approach to create a set of mutations on the fixed version of the program. If the test cases that kill a mutation also fail on the real bug, we consider that the mutation generation approach has successfully “detected” this real bug. To quantify this effectiveness, we compute the Real Bug Detection Rate, which represents the proportion of real bugs for which at least one generated mutation is detected by the same test cases that reveal the real bug. This metric provides insight into how well LLM-generated mutations align with actual software defects in a testing context.
RQ1.2: Average Ochiai Coefficient:  This metric estimates the semantic similarity between a mutation and its corresponding real bug based on their passing and failing test cases (Offutt and Untch, 2001). Ochiai coefficient measures how a mutation resembles its corresponding real bug. The Ochiai coefficient was originally introduced as a general similarity measure but later became widely adopted in spectrum-based fault localization (SBFL) (Zou et al., 2019). In this study, we use the Ochiai coefficient in its original form to assess how closely a mutation resembles its corresponding real bug, which is consistent with the study of Ojdanic et al. (Ojdanic et al., 2023).
More concretely, let $m$ and $b$ be a mutation and its corresponding real bug, and let $fT_{m}$ and $fT_{b}$ be the set of their failing test cases respectively. The coefficient of $m$ is computed by the following formula:
(1) $Ochiai(m,b)=\frac{|fT_{m}\cap fT_{b}|}{\sqrt{|fT_{m}|\times|fT_{b}|}}$
Therefore, if the Ochiai coefficient of $m$ is $1$, it is semantically equal to its corresponding bug $b$. Given a bug $b$, its Ochiai coefficient is calculated as the mean Ochiai coefficient across all its corresponding de-duplicated mutations (i.e., $(C_{b}-D_{b})$), shown as the following formula:
(2) $Ochiai_{b}=\frac{\sum_{m\in(C_{b}-D_{b})}Ochiai(m,b)}{|C_{b}-D_{b}|}$
Given a mutation generation approach and a set of bugs $B$, the Average Ochiai Coefficient of the approach is calculated as the mean Ochiai coefficient across all bugs, shown as the following formula:
(3) $AOC=\frac{\sum_{b\in B}{Ochiai_{b}}}{|B|}$
RQ1.2: High Ochiai Bug Count:  Following the study (Ojdanic et al., 2023), a bug $b$ and a set of mutations are considered to have high Ochiai if they satisfy $(Ochiai_{b}\geq 0.8)$. This threshold indicates a strong semantic similarity between the generated mutations and the real bug. For each mutation generation approach, we compute the number of bugs that meet this criterion.
RQ1.2: Coupling Rate:  Given a real bug, one mutant is considered coupling if the set of test cases that kill the mutant overlaps with the bug-triggering tests. Note that each bug we analyzed has at least one failing test to reveal the bug, called bug-triggering test. The Coupling Rate refers to the proportion of coupled mutants within the set of de-duplicated mutations (i.e., $(C-D)$). Note that Real Bug Detection Rate estimates the effectiveness via the proportion of detected bugs, which Coupling Rate estimates it via the proportion of mutations.
RQ1.2: Coupling Categorization:  We categorize each generated mutation into distinct coupling scales based on their similarity to real bugs, following the classification proposed by (Gay and Salahirad, 2023), and analyze their distribution across different approaches. Specifically, given a compilable mutation (i.e., an element from the set $C$) on the fixed version of a real bug (which can be uncovered by at least one triggering test), we can classify it into the following six scales based on its behavior:
  • Strong Substitution: All triggering tests fail and no additional tests fail, which represents the mutation is a semantic replacement of the real bug under the given test suite.
  • Strong Substitution + Additional Tests Fail: All triggering tests fail but additional non-triggering tests fail.
  • Partial Substitution: Some triggering tests fail and no additional tests fail.
  • Partial Substitution + Additional Tests Fail: Some triggering tests fail and some additional non-triggering tests fail.
  • Not Substitution: Only non-triggering tests fail.
  • Not Detected: No tests fail (i.e., the mutation is alive under the given test suite).
Later, we will classify the mutations of each approach and analyze their distributions.
RQ1.3: AST Node Diversity:  We prefer the set of mutations with diverse syntactic forms, i.e., method invocations, field accesses, etc. Therefore, we parse the code before and after mutation, and check the diversity of newly introduced AST nodes. The approach with a greater variety of AST node types is considered superior, as it indicates a broader exploration of the syntactic space and enhances the likelihood of detecting weaknesses in the test suite.
RQ1.3: AST Edit Distance:  The AST edit distance measures the structural difference between the original and mutated code by quantifying the number of tree transformation operations (i.e., insertions, deletions, and modifications) needed to convert one AST into another. This metric evaluates the extent of syntactic changes introduced by mutations. A higher edit distance suggests more significant syntactic changes.
2.4.2. RQ2: Validity
We consider three key metrics to evaluate the validity of different mutation generation approaches: generation efficiency, compilability, and the proportion of duplicate and equivalent mutations, providing a basic understanding of the validity of each approach.
RQ2.1: Compilability Rate:  the proportion of mutations that successfully compile, is computed as:
(4) $CR=\frac{\enspace|C|\enspace}{|A|}$
Because LLMs cannot guarantee the generation of entirely correct code and the compilation process can be time-consuming, we prefer the approach with a higher comparability rate.
RQ2.2: Duplicate Mutation Rate:  the proportion of duplicate mutations, computed as:
(5) $DMR=\frac{\enspace|D|\enspace}{|A|}$
LLMs sometimes generate duplicate mutations, i.e., mutations that are exactly the same as the original code. While traditional approaches rarely generate duplicate mutations, LLMs can produce them due to their probabilistic text-generation mechanisms and learned patterns from large-scale training data.
RQ2.3: Equivalent Mutation Rate:  the proportion of equivalent mutation, computed as:
(6) $EMR=\frac{\enspace|E|\enspace}{|C-D|}$
Equivalent mutations are semantically equivalent to the original program, which impacts the accuracy of the mutation score. We prefer the approach with a lower EMR.
2.4.3. RQ3: Efficiency
Efficiency describes how fast a mutation approach produces mutations. Mutation testing requires numerous mutations to evaluate the robustness of a test suite effectively. Therefore, the efficiency of mutation generation directly impacts the feasibility and scalability of mutation testing. To assess this, we compare these approaches based on time costs.
RQ3: Average Generation Time:  measures the average time required to generate each mutation, providing a quantitative evaluation of the efficiency of each approach.
2.5. Mutation Generation
In this subsection, we discuss the setups for mutation generation approaches, especially the setups for LLM-based generation. We performed our experimental analysis on four rented cloud servers, each equipped with 2 NVIDIA GeForce RTX 3090 Ti, 168GB memory, and a 64 Core CPU Intel(R) Xeon(R) Platinum 8350C.
2.5.1. Mutation Generation via Traditional Approaches
For the well-established mutation generation approaches (i.e., PIT, Major, LEAM, and $\mu$Bert), we activate all their mutation operators. Since we try to compare the similarity between the generated mutations and real bugs, we filter the mutations within the context of real-world bugs. Note that for all mutation generation approaches, we employ the same setting for the code context.
2.5.2. Mutation Generation via LLMs
Recently, LLMs have shown great potential in numerous code-related tasks, to better understand their ability in mutation testing, we also study LLMs for generating mutations.
@@ 2.5.2. Mutation Generation via LLMs
LLMs operate based on prompts, which consist of natural language instructions, often accompanied by code snippets. LLMs are driven by prompts, which are natural language instructions maybe with code. To effectively leverage LLMs for mutation generation, we design a well-crafted prompt template, shown as Figure 1. We follow the best practices suggesting prompts should comprise four aspects, namely, Instruction, Context, Input Data, and Output Indicator (Guo et al., 2024; Ma et al., 2024). In the Instructions, we direct the LLMs to generate mutants for the target code element. In the Context, we clearly state that mutant is the concept of mutation testing, and additionally provide various sources of information to observe the impact on the performance of LLMs, including the whole Java method surrounding the target code element and few-shot examples sampled from real-world bugs from another benchmark. In the Input Data, we specify the code element to be mutated and the number of mutants to generate. In the Output Indicator, we specify the JSON file format for mutation outputs, facilitating further experiments. In addition, our prompts need few-shot examples, which should originate from real bugs enabling LLMs to learn how to mutate code from real bugs. This allows the model to understand the context and logic behind the changes, improving its ability to generate relevant and effective mutations.
@@
To avoid the few-shot examples leaking information, beyond Defects4J and ConDefects, we employ another benchmark QuixBugs (Lin et al., 2017), which comprises 40 real-world Java bugs and is commonly used by the testing and debugging community (Zhu et al., 2023; Jiang et al., 2023; Xia and Zhang, 2022; Li et al., 2023b). Referring to the study by Deng et al. (Deng et al., 2024), LLMs perform best in generating valid code when using around six few-shot examples. Therefore, we sample six bugs from QuixBugs for few-shot learning in our prompt. To guarantee the diversity of the examples, we randomly select one bug from the dataset and assess whether its modification pattern is similar to the examples already collected. If it is different, we add it to our collection and continue, until we have collected all six examples, shown as Table 3.
To determine the number of mutations generated in a prompt, we adopt the setting of PIT (Coles et al., 2016). We use PIT-generated mutations with the universal set of operators and find that PIT generates 1.15 mutations per line of code. Therefore, based on the given code context, we generate one mutation per line. We explore the influence of context length on mutation generation in Section 4. Additionally, for the temperature, we use the default value for each LLM.
Figure 1. The Prompt Template for LLMs